Find the slope and y intercept for the equation
x+15y=1

Answer by ReadingBoosters(3246)

y = mx+b as linear equation
Using
x+15y=1
15y = 1 - x
y = 1/15 - x/15
slope = -1/15
y intercept = 1/15